OpenStudy (anonymous):

Evaluate: [(40 + 5) − 32] ÷ 9 • 2

OpenStudy (anonymous):

order of operations also known as PEMDAS

OpenStudy (goldphenoix):

You do order of operation. So you do the parenthesis first. 40 + 5 = 45 [45 - 32] / 9 * 2 Do the other parenthesis (bracket) 45-32 = 13 13 / 9 * 2 Now do the division. 13 / 9 = 1.44. Final step is to do the multiplication. 1.44 * 2 = 2.88. I hope this help! :)